Paint Town sold 45 paintbrushes, one kind at $8.50 each and another at $9.75 each. In all, $398.75 was taken in for the brushes. How many of each kind were sold?
Answer by Maths68(1474)

Let
one kind = x
other kind = y
Total paintbrushes sold
x+y=45
x=45-y ……………..(1)
Total amount received
8.5(x)+9.75(y)=398.75 …………(2)
Put the value of x in (1)
8.5(45-y)+9.75y=398.75
382.50-8.5y+9.75y=398.75
-8.5y+9.75y=398.75-382.50
1.25y=16.25
y=16.25/1.25
y=13
Put the value of y in
x=45-13
x=32
one kind = 32 brushes sold
other kind = 13 brushes sold